LFBG Re-Commissioning with LUCI: 16 September 2010 UT

Observer: DThompson, JHill (LBTO)
Telescope Operator: SAllanson and GBechetti (LBTO)
Telescope Support: JUrban (LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Seeing wasn't so great tonight (0.7 - 1.5 arcsec), perhaps because we had more solar heating during the day. We updated the collimation models and the pointing model for LFBG. We checked the transformation measured last night and analyzed by Jesper. After getting the flips and rotations straightened out, the residuals were 33/31 microns over a small range of position angles. When we rotate over all position angles in a given place on the sky, the residuals (measured on the LUCI science detector) are dominated by LUCI internal flexure (even with flexure compensation on). We tested automatic field aberration data collection scripts in IRAF and IDL. We checked the DETXY offsets with the new transformation. Seeing improved to 0.5 - 0.8 arcsec at the end of the night, so we got some decent field data with the -20 arcsec Tilt correction on M2 from last night.

3-Night Overview

We've nearly finished the important AGw3 and LUCI recommissioning activities. Just a few tasks remain to be done:
  • Refine pupil position parameters for WFSing in IDL and GCS (Doug off-sky), and install those.
  • Complete the Field Aberration Measurements and M2 alignment (Andrew, et al.). This needs about 0.5 night of sky time in decent seeing. Andrew says more time could be needed to sort out the M3 angle. DONE (21-SEP)
  • Adjust the zeropoint of the new pointing model after the Field Aberration adjustments are completed. DONE (22-SEP)
  • Adjust the global offsets in the M2 collimation model to include the Field Aberration adjustments. DONE (22-SEP)
  • We have not explicitly measured the AGw-to-LUCI focus offset, but it seems to be very close to optimal as we had images at 0.3 arcsec FWHM on LUCI.
  • We have not re-measured the photometric or astrometric performance of the guider CCD, but that is not very important. Nor do we have reason to think it will have changed.

See Com20100921 for the next round of field aberration measurements

Preparations

Tony has patched Telemetry r10 to fix Issue #2832.

Jesper has created a new transformation based on last night's measurements. We have installed it in oacontrol.conf.

Added these global offsets from last night to the SXPMLUCIFERCollimation.dat file.
Xglobal=-0.684 Yglobal =2.196 Zglobal=+0.186 RXglobal=+23.73 RYglobal=+6.83
These globals were only used for collecting the collimation model data until 04:00 UT.

Details

Open

01:40 Open T=12.2 degC, D=2.4 degC, Wind light from west. A few scattered low clouds.

01:46 Dave adjusts the pointing on a bright star. IE=-58.4 CA=-23.4 New collimation adjustment a big improvement.

01:53 Preset ACTIVE BS9178 PA=180. Dense field, take astrom data. Luci#0006, 0007=sky Ks+Br_gam 6x10.0s.

Update AGw transform with Jesper's new solution from last night's data.

02:03 Preset again to BS9178 PA=180. Luci#0008, 0009; 0010 at PA=150

Check current collimation model.

Collimate on BS9178 EL=52
Check on ACT0081 EL=20 not good
Check on ACT0302 EL=83 OK

Take collimation model data.

Star EL NNNN Notes
BS9170 52.5 0013 parallel offset Z -1.0 mm after this image
BS9160 34.1 0014
BS9155 21.1 0015 Poor collim, elongated, shading across pupil?
BS9160 32.5 0016
BS9164 44.1 0017
BS9170 49.0 0018
BS9183 55.6 0019 correct the pointing, then...
FS149 81.7 0022
FS151 69.8 0023
FS152 49.0 0024
BS9169 56.5 0025
BS9177 69.8 0026
BS9183 60.3 0027
BS9182 50.7 0028 parallel offset Y -2.0 mm after this image
FS149 87.2 0029
FS151 78.3 0030

Header data for these images collected in /home/lbcobs/Collimation/new/LUCI_20100916.txt. That file has already been corrected for the parallel offsets of M1/M2 implemented above.

Analyzed this data with /home/lbcobs/Collimation/new/mkcoll_irtc.pro. See attached plot of the results below.

Take pointing model data.

pointing corrected IE=-73 CA=-26 after implementing new collimation model SXPMLUCIFERCollimation.20100916b.dat and SXSMLUCIFERCollimation.20100916b.dat

mount log file started: /home/telescope/20100916.ptlog
04:23 Preset to BS9183 to collimate

P5 UT SourceID LUCI PM-Corrected Coordinates Notes
01 04:29 WT10_373 0033 22:25:52.5870 53:56:40.714  
02 04:33 ACT0407 0034 5:00:53.9876 77:46:31.820  
03 04:35 WT10_401 0035 15:37:45.7965 71:09:13.378  
04 04:36 WT10_358 0036 18:56:32.1475 46:45:58.154  
05 04:38 ACT0281 0037 15:44:55.8963 21:26:41.171  
06 04:40 WT10_179 0038 18:24:30.0223 -8:10:30.615 unwrap AZ
07 04:45 ACT0253 0039 20:09:00.7119 13:00:57.024  
08 04:48 WT10_068 0040 19:57:37.1584 -37:44:41.772  
09 05:02 WT10_168 0041 22:54:07.1594 -12:11:26.003 AZ glitch @4:51, IT#2836
10 05:06 WT10_274 0042 22:32:27.4868 19:37:25.416  
11 05:08 WT10_225 0043 1:24:54.5826 3:20:57.709  
12 05:10 ACT0325 0044 1:09:17.3673 34:48:21.132  
13 05:12 WT10_338 0045 0:19:46.6816 36:57:44.574  
14 05:13 WT10_278 0046 1:22:24.7100 22:05:11.843  
15 05:15 WT10_242 0047 0:52:28.8808 13:38:55.429  
16 05:16 ACT0306 0048 22:11:23.1754 29:15:16.449  
17 05:18 ACT0185 0049 22:32:34.2678 -6:28:01.540  
18 05:20 ACT0085 0050 22:22:04.5978 -34:29:19.770  
19 05:22 WT10_115 0051 21:31:51.3537 -25:24:45.856  
20 05:23 ACT0289 0052 21:35:34.2115 23:58:39.601  
21 05:26 WT10_270 0053 19:34:45.3084 16:38:29.937  
22 05:28 ACT0213 0054 17:21:06.9382 1:26:37.913  
23 05:30 WT10_300 0055 17:52:32.1803 27:11:36.962  
24 05:31 WT10_334 0056 20:51:30.9327 37:59:22.148  
25 05:34 ACT0398 0057 21:09:31.6912 68:29:26.728  
26 05:36 ACT0406 0058 14:18:36.9088 75:40:07.535  
27 05:37 WT10_410 0059 1:55:42.7541 81:57:32.080  
28 05:39 ACT0321 0060 21:56:08.7664 34:33:01.307  
01 05:42 WT10_362 0061 22:51:50.8461 46:44:07.784 repeat first 12 pos'n
02 05:44 WT10_407 0062 6:35:26.9137 80:34:23.631  
03 05:46 ACT0396 0063 17:33:18.0427 67:12:13.905  
04 05:48 ACT0347 0064 20:13:23.7982 43:48:40.236  
05 05:49 ACT0283 0065 17:09:27.2491 22:05:29.951  
06 05:52 WT10_181 0066 19:40:45.1713 -7:30:45.061  
07 05:54 WT10_255 0067 21:32:59.1655 12:32:22.292  
08 05:55 WT10_070 0068 21:29:49.6151 -37:41:34.184  
09 05:58 ACT0170 0069 0:27:06.9215 -9:52:23.220  
11 06:00 WT10_227 0070 2:41:21.6144 4:25:39.460 (cell 10 empty)
12 06:01 ACT0327 0071 2:46:06.7721 36:16:00.305  
10 06:03 WT10_276 0072 23:58:19.1232 19:41:36.791  

Pointing model plot from TPOINT:
tpmodel.jpg

Sky RMS = 4.87 arcsec. File LUCIFERSX.20100916.ptmod copied to the usual place and made the default in PCSInstrument.conf. PCS restarted ~UT7:40 to read in the new model.

Check the New Transformation

06:17 pointto StoneOagw gs=20 PA=180 Tel="ACTIVE" 0074 1036.4 1046.0 # EL=55
06:22 pointto StoneOagw gs=12 PA=180 Tel="ACTIVE" 0075 1036.4 1033.5
06:25 pointto StoneOagw gs=14 PA=180 Tel="ACTIVE" 0076 1028.3 1012.1

Turn on LUCIFER Flexure compensation and re-check
06:28 gs=14 PA=180 0077 1028.2 1013.1
06:29 pointto StoneOagw gs=12 PA=180 Tel="ACTIVE" 0078 1036.6 1033.6
offset to PA=225 0079 1028.0 1039.9
offset to PA=135 0080 1037.6 1020.5
06:38 pointto StoneOagw gs=20 PA=180 Tel="ACTIVE" 0081 1036.5 1046.0 # EL=53

Replace previous transformation:
06:44 pointto StoneOagw gs=20 PA=180 Tel="ACTIVE" 0082 1027.7 1046.4 # EL=52
06:46 pointto StoneOagw gs=14 PA=180 Tel="ACTIVE" 0083 1023.1 1029.0

06:52 pointto StoneOagw gs=3 PA=180 Tel="ACTIVE" 0084 1028.6 1035.4
06:54 pointto StoneOagw gs=3 PA=225 Tel="ACTIVE" 0085 1029.2 1029.7

06:57 pointto StoneOagw gs=7 PA=270 Tel="ACTIVE" 0086 1028.6 1035.4
07:00 pointto StoneOagw gs=9 PA= 90 Tel="ACTIVE" 0087 1031.2 1028.7

Neither transformation is very good!
For now, AGw rotation center at ~~ 1029.90 1032.05 on LUCIFER (use for pointing model)

Script Development

07:02 Stop GCS WFSing on the present star

after some fooling around with output redirection and fixing typos, I've got agw_collimate.py working correctly so it can call Doug's script collimate_multiple.pro.

Practice Field Aberration Data on BS9107

This is mainly practice data as the seeing is worse than 1 arcsec.

07:52 Apply -20 arcsec of Tilt to M2 (Andrew and Doug have a complicated debate about this.) This gives adjustments of X=+1.021 RY=196.40 to M2.

working in /OldRepository/AGW_Data/100916/
This first run has the interactive threshold set to a large number (400) for testing.

08:01 cl < home$field_BS9107c.cl

GS wfscimage
0 5-8
82 9-11
0 12-13
35 14-16
0 17-18
19 19-21
0 22-23
64 24-26
0 27
96 28-30
0 31-34

manually interrupted here........

Testing Doug's Field Data script

08:35 Doug starts testing his script on BS9107 - seeing is 0.72 - 0.95 arcsec.

He's working in the parallel directory /OldRepository/AGW_Data/20100916/
Note that once again Doug and John have parallel directories 20100916 and 100916. Doug's numbering of the wfsc images starts again at 0001.

08:40 Probe couldn't reach gs=70 (that it could previously reach).
08:49 Probe couldn't reach gs=68 y=262 - This is because he was accidentally using a different central star as the target.

09:09 retesting with the normal central star. Now we have the same star, but Doug(IDL) and Dave(IRAF) don't agree on the proper motion corrected coordinates of this star (gs=0 = gs61).

Note that the IDL pupil parameters need fine tuning when the guide star is far off-axis. Doug will need to fix that when he analyzes this data. Then we will propagate it to AIP_L.cfg.

09:28 now John running field_collect as LBTO. Seeing 1 arcsec. Doug goes to bed.

GS wfscimage
0 19-24
82 25-27
0 28-30
35 31-33
0 34-39

09:47 didn't converge in 6 iterations, so we stop and do something else.

Another Transformation Check

Jesper thinks he may have gotten his X-flips and Y-flips confused. So he's now sent us a new transformation in oacontrol.conf.20100915new2

09:48 Installed the new transformation file

09:50 Preset to ...... and try different guide stars

The good news is that oacontrol.conf.20100915new2 is a much better transformation. So whatever Jesper flipped was the right thing.

Jesper reports:
... I got tangled in the usual flipping and
rotation circus but now things are clear and I believe I understand exactly
what is going on. I have attached my little awk-script for preparing the data
for the fitting program so you can include the LUCIFER variant in your IDL
scripts. I suppose that using LUCIFER will be the standard calibration
procedure in the future, so I thought I might spare you some flipping and
rotating...

  I should note that the last fit had stddev values of (x,y)=(40,30)microns.
If I reject one of the 22 stars (possible proper motion error?) I get a stddev
of (33,31), so this should really be as good as it gets. At this level it
seems that the LUCIFER flexure contributes significantly to the errors so to
reach this low stddev the PA should be kept within a small range.

DETXY Offsets (check offset orientation along slit)

10:12

Script offset NNNN Xcen___Ycen
OFFSET = 0.0 60.0 0103 1012.9 1541.8
OFFSET = 0.0 -120.0 0104 1013.2 544.1
OFFSET = 0.0 120.0 0105 1012.9 1541.8
OFFSET = 0.0 -120.0 0106 1013.3 543.9
OFFSET = 0.0 120.0 0107 1012.9 1541.8
OFFSET = 0.0 -120.0 0108 1013.5 543.6
OFFSET = 0.0 120.0 0109 1013.0 1541.4
OFFSET = 0.0 -120.0 0110 1013.4 543.7
OFFSET = 0.0 120.0 0111 1012.9 1541.3

Had problems with NaNs in the Rotator Polys. See Issue 2839. Stopping and restarting PCS got rid of them.

Another shot at BS9107 Field Aberration Data

10:56 field_collect with bs9107_field.list - seeing is 0.5 - 0.9 arcsec on guider

GS wfscimage
0 40-44
82 45-47
0 48-50
35 51-53
0 54-55
19 56-58
0 59
64 60-62
0 63
96 64-66
0 67-68
40 69-71
0 72-73
35 74-76
0 77
48 78-80
0 81
8 82-84
0 85
68 86-88
0 89-90
96 91-93
0 94-95
39 96-98
0 99
97 100-102
0 103

12:03 field_collect with bs9107_field.list - seeing is 0.5 - 0.7 arcsec on guider

GS wfscimage
0 104-105
82 106-108
0 109
35 110-112
0 113-114
19 115-117
0 118
64 119-121
0 122

Andrew reports the next afternoon that we got the magnitude and the axis of the M2 pointing adjustment right, but the sign was wrong. We should have applied +20 Tilt to M2 (see below).

Note that IDL was sending Zernikes Z4 through Z20 for this dataset. See Com20100921.

Close

12:26 Close for impending dawn, T=11 degC, D=1 degC, Wind light from West. Sky is clear.

After observing, I changed WFS_rmsLimit was 600.0 now 800.0 in AIP_L.cfg for AO5. We've verified in 0.45 arcsec seeing that after you've sent a correction of 800 nm, the image is good (below 0.5 arcsec). I even observed a 0.5 arcsec image after a 2700 nm correction, but let's not get carried away.

On the afternoon of 17-SEP-2010 UT, I also added the +20 Tilt to M2 in a permanent way in the global offsets. The previous global offsets (also M2 total collimation) were: X=-7.010 Y=-2.600 Z=-1.000 RX=148.650 RY=181.800, and the new values are X=-8.031 Y=-2.600 Z=-1.000 RX=148.650 RY=-14.599. This is pushing us far out on the X-range of M2, so a swing arm adjustment may be needed if the next round of focal plane tilt measurements confirm this position. Note also that the Y-value of -2.6 was influenced by a parallel offset to M1 and M2 to get collimation range in warm weather.
See the revised values in Com20100921.

-- JohnHill - 15 Sep 2010

Topic attachments
I Attachment Action Size Date Who Comment
AIP_L.cfgcfg AIP_L.cfg manage 9 K 16 Sep 2010 - 10:01 JohnHill GCS configuration file for AGw3
LUCI_20100916.gifgif LUCI_20100916.gif manage 25 K 16 Sep 2010 - 04:25 JohnHill M1 Collimation Model Plot with temperature contours
SXPMLUCIFERCollimation.20100916b.datdat SXPMLUCIFERCollimation.20100916b.dat manage 1 K 16 Sep 2010 - 04:23 JohnHill M1 Collimation Model
SXSMLUCIFERCollimation.20100916b.datdat SXSMLUCIFERCollimation.20100916b.dat manage 1 K 16 Sep 2010 - 04:24 JohnHill M2 Collimation Model
daq.20100916.jpgjpg daq.20100916.jpg manage 691 K 22 Sep 2010 - 03:34 JohnHill Plot of Telescope Temperatures
tpmodel.jpgjpg tpmodel.jpg manage 116 K 16 Sep 2010 - 07:43 DavidThompson Pointing model residuals
Topic revision: r20 - 22 Sep 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback